# Volumetric Semantically Consistent 3D Panoptic Mapping

### v1

@@ III-D Instance Refinement
where $\mathcal{E}_{C}(O^{C}_{i})$ is the set of edges representing semantic class $C$ of the superpoints in $O^{C}_{i}$.
We iterate over nodes $L_{S}\in\mathcal{V}_{C}$, and detach node $L_{S}$ from its initial guess $O_{L_{S}}$ if it has low edge confidence within $O_{L_{S}}$: $P^{C}_{L_{S}}(O_{L_{S}})<\theta^{C}_{D}P_{O}(O_{L_{S}})$, where $P^{C}_{L}(O)$ is the largest edge confidence ${P_{\epsilon}}^{C}(L,L_{S})$ between superpoint $L$ and other superpoints $L_{S}$ stemming from instance $O$. We select the node with the highest confidence $P_{L}^{C}(L_{S})$ among the detached nodes and, progressively, the less confident ones. For each detached node $L_{S}$, we iterate through its neighbors and assign $O_{\text{neigh}}$ to it, if $P^{C}_{L_{S}}(O_{\text{neigh}})>\theta^{C}_{O}P_{O}(O_{\text{neigh}})$ and $P^{C}_{L_{S}}(O_{\text{neigh}})>\theta^{C}_{L}P_{L}^{C}(L_{S})$. If no such neighbors exist, $L_{S}$ will be assigned to a new instance label $O_{\text{new}}$. Neighbors of node $L_{S}$ are defined as nodes connected with $L_{S}$ in a spatial-panoptic manner: $P_{\epsilon}^{C}(L_{S},L_{\text{neigh}})>0$.
@@ IV EXPERIMENTS
We evaluate on the SceneNN [25] and ScanNet v2 datasets [28], containing RGB-D sequences, ground truth (GT) 3D meshes, camera poses, and semantic instance annotations. We compare the proposed method with multiple state-of-the-art frameworks [20, 21, 1, 26, 27, 17]. We use four metrics to compare accuracy, computed by thresholding the intersection over union (IoU) at thresholds $0.5$ and $0.75$: (i) mean average precision (mAP), (ii) total number of true positive predicted instances ($N_{\text{TP}}$), (iii) panoptic quality (PQ) [15], and the (iv) Intersection over Union between GT instance and segmented superpoints ($\text{IoU}_{L_{S}}$), defined as $\text{IoU}_{L_{S}}=\frac{1}{\left|O\right|}\sum_{o\in O}\sum_{L_{S}}\text{IoU}(L_{S},o)$, where $O$ represents the set of all GT instances. Metric $\text{IoU}_{L_{S}}$ measures the segmentation accuracy of superpoints, and is only for methods performing superpoint segmentation and having public source code [1, 21]. We use the [23] panoptic segmentation network for all methods.
@@
The pipeline is implemented in C++ and runs on an Intel-12700K CPU. The RGB panoptic segmentation runs on an Nvidia RTX3090. We run [1, 21] on same machine, while [17] is run on an E5-2680 CPU and Titan Xp GPU due to library dependency issue. For all sequences in SceneNN and ScanNet v2, we use the same hyper-parameters: $\theta_{merge}=3$ (Sec. III-B), $K_{C}=15.0,\theta=0.5$ (Sec. III-C), Parameters $\theta^{C}_{O}$, $\theta^{C}_{D}$, $\theta^{C}_{L}$ (Sec. III-D) are tuned on ScanNet for 18 semantic categories. The Label-TSDF voxel size is $1$cm.
@@ IV-C Ablation Study
Table II(b) displays the ablation study with components removed from the proposed method. Label ”w/o III-B” means we use the superpoint segmentation module from [21] instead of the proposed one; ”w/o III-C and w/o III-D” means no semantic regularization and no instance refinement, respectively; ”w/ GT Mask” means we use ground truth panoptic masks in III-A by rendering from GT meshes instead of using predictions from [23]. We show results both in the GT pose and SLAM cases. While the two cases have a significant performance gap, ranking the components by importance seems similar. The conclusions are as follows:
@@
    Method 1 compared to 3, 1 to 4 and 1 to 7. The results show that the methods from both Sections III-B and III-D play significant roles in improving the accuracy of semantic-instance segmentation.
@@ IV-D Run-time and Memory Analysis
Table IV(a) shows the run-time and memory analysis of each component in the pipeline. The 2D panoptic-geometric segmentation and fusion in Section III-A takes the majority of the time and memory usage. It is important to note that this largely depends on the employed panoptic segmentation network. Thus, changing the network substantially changes the required time and memory. With limited computational resources, a lightweight panoptic segmentation (e.g., [29]) and lower voxel resolution can be used.
@@ V CONCLUSIONS
In this paper, we introduce a novel 2D-to-3D semantic instance segmentation algorithm. Drawing upon a Voxel-TSDF representation, our method combines panoptic prediction confidence, semantically consistent superpoints, and graph-optimized semantic labeling and instance refinement. These components allow the proposed algorithm to outperform current state-of-the-art techniques on public and large-scale datasets. We achieve this superior accuracy while maintaining real-time processing speeds comparable to less accurate alternatives. Although the proposed approach was inspired by Voxblox++, our results surpass its accuracy, on average, by a significant 16.1 mAP points, underscoring the impact of our improvements. Furthermore, we draw attention to the inherent pitfalls of the conventional practice of relying on ground truth camera poses as input as it introduces considerable disparities between reported outcomes and real-world performance.

### v2

@@ III-D Instance Segmentation via Super-Point Graph
where $\mathcal{E}_{C}(O^{C}_{i})$ denotes the set of edges connecting super-points assigned to $O^{C}_{i}$. Super-points $L_{S}\in\mathcal{V}_{C}$ are then evaluated to determine if they should remain within their initial instance grouping $O^{C}_{i}$ based on the strength of their connections to other super-points within the same instance. Those with insufficiently strong connections (below a threshold $\theta^{C}_{D}P_{O}(O^{C}_{i})$) are considered for reassignment.
For each super-point $L_{S}$ undergoing reevaluation, we explore its neighbors to identify potential reassignments to instances with strong connections, satisfying conditions $P^{C}_{L_{S}}(O_{\text{neigh}})>\theta^{C}_{O}P_{O}(O_{\text{neigh}})$ and $P^{C}_{L_{S}}(O_{\text{neigh}})>\theta^{C}_{L}P_{L}^{C}(L_{S})$. Neighbors are defined based on spatial-panoptic connectivity with $P_{\epsilon}^{C}(L_{S},L_{\text{neigh}})>0$. Super-points without suitable neighbors are allocated a new instance label $O_{\text{new}}$.
@@ IV EXPERIMENTS
We conduct experiments and compare the proposed method with multiple state-of-the-art frameworks [21, 2, 1, 26, 27, 3] on the SceneNN [25] and ScanNet v2 datasets [28], containing RGB-D sequences, ground truth (GT) 3D meshes, camera poses, and semantic instance annotations. We use four metrics to compare accuracy, computed by thresholding the intersection over union (IoU) at thresholds $0.5$ and $0.75$: (i) mean average precision (mAP), (ii) total number of true positive predicted instances ($N_{\text{TP}}$), (iii) panoptic quality (PQ) [17], and the (iv) Intersection over Union between GT instance and segmented super-points ($\text{IoU}_{L_{S}}$), defined as $\text{IoU}_{L_{S}}=\frac{1}{\left|O\right|}\sum_{o\in O}\sum_{L_{S}}\text{IoU}(L_{S},o)$, where $O$ represents the set of all GT instances. Metric $\text{IoU}_{L_{S}}$ measures the segmentation accuracy of super-points, and is only for methods performing super-point segmentation and having public source code [1, 2]. We use the [23] panoptic segmentation network for all methods.
@@
The pipeline is implemented in C++ and runs on an Intel-12700K CPU. The RGB panoptic segmentation runs on an Nvidia RTX3090. We run [1, 2] on same machine, while [3] is run on an E5-2680 CPU and Titan Xp GPU due to library dependency issue. For all sequences in SceneNN and ScanNet v2, we use the same hyper-parameters: $\theta_{merge}=3$ (Sec. III-B), $K_{C}=15.0,\theta=0.5$ (Sec. III-C), Parameters $\theta^{C}_{O}$, $\theta^{C}_{D}$, $\theta^{C}_{L}$ (Sec. III-D) are tuned on ScanNet for 18 semantic categories. The Voxel-TSDF voxel size is $1$cm.
@@ IV-C Ablation Study
Table II(b) displays the ablation study with components removed from the proposed method. Label ”w/o III-B” means we use the super-point segmentation module from [2] instead of the proposed one; ”w/o III-C and w/o III-D” means no semantic regularization and no instance refinement, respectively; ”w/ GT Mask” means we use ground truth panoptic masks in III-A by rendering from GT meshes instead of using predictions from [23]. We show results both in the GT pose and SLAM cases. While the two cases have a significant performance gap, ranking the components by importance seems similar. The conclusions are as follows:
@@
    Method 1 compared to 3, 1 to 4 and 1 to 7. The results show that the methods from both Sections III-C and III-D play significant roles in improving the accuracy of semantic-instance segmentation, qualitative results are shown in Fig. 6.
@@ IV-D Run-time and Memory Analysis
Table. IV(a) shows the run-time and memory analysis of each component in the pipeline. The 2D panoptic-geometric segmentation and fusion in Section III-A takes the majority of the time and memory usage. It is important to note that this largely depends on the employed panoptic segmentation network. Thus, changing the network substantially changes the required time and memory. With limited computational resources, a lightweight panoptic segmentation (e.g., [29]) and lower voxel resolution can be used.
@@ V CONCLUSIONS
In this paper, we introduce a novel 2D-to-3D semantic instance segmentation algorithm. Drawing upon a Voxel-TSDF representation, our method combines panoptic prediction confidence, semantically consistent super-points, and graph-optimized semantic labeling and instance refinement. These components allow the proposed algorithm to outperform current state-of-the-art techniques on public and large-scale datasets. We achieve this superior accuracy while maintaining real-time processing speeds comparable to less accurate alternatives. Although the proposed approach was inspired by Voxblox++, our results surpass its accuracy, on average, by a significant 16.1 mAP points, underscoring the impact of our improvements. Furthermore, we draw attention to the inherent pitfalls of the conventional practice of relying on ground truth camera poses as input as it introduces considerable disparities between reported outcomes and real-world performance. We will release the source code.